Suppose the government of a small nation has provided the following data:

Adult Population 692,000
Full-time Workers 401,360
Part-time Workers 14,532
Unemployed workers actively looking for work 36,361
Unemployed workers who have become discouraged 11,764
Part-time workers under 16 2,076
Unemployed workers under 16 actively looking for work 692

Instructions: Round your answers 2 decimal places.

What is the labor force participation rate? ____ percent

What is the unemployment rate? _____ percent
Answer #1

Labor participation ratio means calculation of labor force which is working,employed or unempolyed and which is having worker above age of 15.hence below 16 age group worker is not included in labor force.working age population means total population.

Unemployment rate means the rate which is calculated number of worker unemployed who is either looking for a job or not .

Labor participation ratio - (labor force ÷ Working age population )* 100

Labor force = 401360+ 14532 +36361 = 452253

Working age population = 692000

Labor participation ratio = (452253 ÷ 692000)*100

= 65.35 %

unemployment rate = (unemployed people ÷ labor force ) * 100

= (36361 ÷ 452253 )* 100

= 8.04 %